OpenStudy (anonymous):

The sum of two consecutive odd integers is 244. What is the smaller integer?

OpenStudy (monica0597):

Do you have any answer choices?

OpenStudy (anonymous):

Take the smaller number as X. X+(X+2)=244 2x+2=244 2x=242 x=121

OpenStudy (anonymous):

119 125 123 121

OpenStudy (anonymous):

so the smaller number is 121

OpenStudy (monica0597):

244-121=123 121 and 123 are consecutive. There u go :)

OpenStudy (anonymous):

ok thank you :)

OpenStudy (monica0597):

U welcome :)

hero (hero):

Message to new users: You should try to guide the student to the answer. This is Open Study, not Open Answer. Giving students a full set of steps without involving them in the process does not guarantee that the user will know how to do the problem on their own. Please try to guide, not give answers. For more info, review the code of conduct: http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

i do understand what i have to do know so there shouldnt be anything wrong ....